# Bayesian Perspective for Orientation Estimation in Cryo-EM and Cryo-ET

### v1

@@ 4 Bayesian rotation determination as part of the volume reconstruction problem
Thus far, we have introduced the MAP and MMSE rotation estimators for estimating the rotation between a single noisy observation $y$ and a reference volume $\overline{V}$, as defined in (2.1). An intriguing question that arises is how these estimators can be utilized and influence performance within a 3D volume reconstruction process, which constitutes the main computational challenge in cryo-EM and cryo-ET.
@@
Structure reconstruction typically follows two primary approaches: hard-assignment or soft-assignment methods, which are generally implemented through iterative refinement. In the hard-assignment approach, each observation is assigned a single orientation based on the highest correlation, and the 3D structure is then reconstructed given the rotations. In contrast, the soft-assignment method assigns probabilities across all possible orientations for each observation, enabling the 3D structure to be recovered as a weighted average of the observations, with weights determined by these probabilities. The iterative application of the soft-assignment procedure aligns with the EM algorithm [12, 46], which serves as the core computational method in modern cryo-EM [44, 40].
@@
In the following, we demonstrate that incorporating the MMSE rotation estimator into the volume reconstruction process in the cryo-ET model (without projections) resembles the expectation-maximization (EM) algorithm, as it accounts for the full distribution of possible outcomes. In contrast, substituting the MAP estimator into the algorithm operates more like a hard-assignment reconstruction method, focusing exclusively on the most likely outcome. This distinction underscores the broader applicability and flexibility of the Bayes estimator in capturing uncertainty and delivering more accurate estimates for structure reconstruction.
@@ 4.1 Connection to the EM algorithm in volume reconstruction without projection
Unlike the previous orientation estimation model (2.1), we consider the following simplified model for volume reconstruction (see also [48, 4, 14]). We observe $M$ i.i.d samples taking the form
$y_{i}=g_{i}^{-1}\cdot V+\varepsilon_{i},\quad\quad i=1,\ldots,M,$ (4.1)
where $V$ is 3D volume structure of interest, $\{g_{i}\}_{i=0}^{M-1}\in\mathsf{SO}(3)$ are unknown latent variables following i.i.d. uniform distribution and $\{\varepsilon_{i}\}_{i=0}^{M-1}$ are i.i.d isotropic Gaussian noise with variance $\sigma^{2}$. The goal is to recover $V$ from the observations $\{y_{i}\}_{i=0}^{M-1}$, treating the rotations $\{g_{i}\}_{i=0}^{M-1}$ as latent variables.
@@
For the case where $V$ represents a 2D image, $\{g_{i}\}_{i=0}^{M-1}\in\mathsf{SO}(2)$ corresponds to in-plane rotations, and the model is used for image recovery (see more in Section 4.2[31].
The most common approach to solve this reconstruction problem is through the EM algorithm that applies soft-assignment iteratively as detailed in Algorithm 2. At each new iteration, the algorithm utilizes the volume estimate from the previous iteration, denoted as $\hat{V}^{(t)}$, to update the estimation of the volume $\hat{V}^{(t+1)}$, based on the observations $\left\{y_{i}\right\}_{i=0}^{M-1}$. The following proposition shows the relation between the update rule of the volume structure estimation at iteration $t+1$, and the MMSE rotation-determination estimator as introduced (2.11).
@@
    Compute for every $10\leq i\leq M-$ 1:
    $p_{i}^{(t)}(g)=\frac{\exp\Big{(}-\left\|y_{i}-\left(g^{-1}\cdot\hat{V}^{(t)}\right)\right\|^{2}/2\sigma^{2}\Big{)}}{\int_{g\in\mathsf{SO}(3)}\exp\Big{(}-\left\|y_{i}-\left(g^{-1}\cdot\hat{V}^{(t)}\right)\right\|^{2}/2\sigma^{2}\Big{)}\mathrm{d}g},$ (4.2)
@@
[…]
    $=$ $\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)(g\cdot y_{i})\mathrm{d}g.$ (4.3)
@@
$\hat{V}^{(t+1)}=\frac{1}{M}\sum_{i=0}^{M-1}\hat{g}_{\mathsf{MMSE},i,t}\cdot y_{i},$ (4.4)
@@
where $\hat{g}_{\mathsf{MMSE},i,t}$ is the rotation MMSE estimator with respect to the uniform prior given observation $y_{i}$ and underlying structure $\hat{V}^{(t)}$ as introduced in (2.11)
$\hat{g}_{\mathsf{MMSE},i,t}=\mathbb{E}[g|y_{i},\hat{V}^{(t)}]=\int_{g\in\mathsf{SO}(3)}\ p_{i}^{(t)}(g)g\mathrm{d}g,$ (4.5)
with $\mathrm{d}g$ corresponds to the uniform distribution over $\mathsf{SO}(3)$.
@@
The proof of this proposition is presented in Appendix C. In words, the proposition shows that the update rule for the volume structure estimation at iteration $t+1$, given the volume $\hat{V}^{(t)}$, is equivalent to aligning each observation $\left\{y_{i}\right\}_{i=0}^{M-1}$ using the associated MMSE estimator $\hat{g}_{\text{MMSE},i,t}$, computed based on the reference volume $\hat{V}^{(t)}$, and then averaging the aligned observations (4.4). Thus, the MMSE rotation-determination estimator is a key ingredient in the EM algorithm for volume reconstruction.
The MAP estimator as part of volume reconstruction.
@@
In contrast to the soft assignment procedure, if we replace the MMSE estimator with the MAP estimator $\hat{g}_{\text{MAP}}$, we obtain the structure reconstruction algorithm by applying hard assignment iteratively. To be more specific, the MAP estimator applied to the observations $\left\{y_{i}\right\}_{i=0}^{M-1}$ can be viewed as a hard assignment among all possible rotations. In practice, this procedure involves making a hard decision where a single rotation is selected from the rotation grid according to the closest alignment. In the $(t+1)$-th iteration, similarly to (4.4), the hard-assignment process can be expressed as follows:
$\hat{V}^{(t+1)}=\frac{1}{M}\sum_{i=0}^{M-1}\hat{g}_{\mathsf{MAP},i,t}\cdot y_{i},$ (4.6)
where $\hat{g}_{\mathsf{MAP},i,t}$ is defined by
$\hat{g}_{\mathsf{MAP},i,t}=$ $\operatorname*{arg\,max}_{g\in\mathsf{SO}(3)}{\mathbb{P}}(g|y_{i},\hat{V}^{(t)})=\operatorname*{arg\,min}_{g\in\mathsf{SO}(3)}\|y_{i}-(g^{-1}\cdot\hat{V}^{(t)})\|^{2}.$ (4.7)
@@ 4.2 Empirical results for volume reconstruction and the “Einstein from Noise” phenomenon
We demonstrate empirically volume reconstruction by applying the MAP estimator $\hat{g}_{\mathsf{MAP}}$ and the MMSE estimator $\hat{g}_{\mathsf{MMSE}}$ as part of the reconstruction problem, as specified in (4.6) and (4.4), respectively.
@@
We demonstrate the reconstruction process, which integrates MAP and MMSE rotation estimation as intermediate steps, using the two examples: 2D image recovery (Figure 4) and 3D volume reconstruction without projection (Figure 5). The iterative reconstruction process, as outlined in Algorithm 2, was performed until convergence (i.e., when the relative difference between consecutive iterations fell below a predefined threshold) or until reaching a maximum of 100 iterations.
@@
Figures 4 and 5 were generated using slightly different methods. For the 2D experiment presented in Figure 4, we used polar coordinates, while in the 3D experiment, shown in Figure 5, we used a standard Cartesian basis. The primary difference lies in the interpolation required for producing Figure 5, which utilizes cubic interpolation for each observation based on the estimated rotation. This introduces certain “quantization” errors. Additionally, as detailed in Section 2, the 3D reconstruction process, presented in Figure 5, requires a “rounding” step which amounts to solving the Orthogonal Procrustes procedure.
In Figure 4, the true and template structures were generated in a polar representation with $d=300$ radial points and $L=30$ polar angle points. The reconstruction process was performed with $M=5\times 10^{4}$ observations. The additive noise was added in the polar representation. In Figure 5, $M=3000$ observations were used, with the rotation group $\mathsf{SO}(3)$ grid size of $L=300$.
@@
Figure 4: Comparison of 2D image recovery using the MMSE and the MAP rotation estimators. Iterative image recovery procedures with the MMSE estimator ($\hat{g}_{\text{MMSE}}$) and MAP estimator ($\hat{g}_{\text{MAP}}$) are defined in (4.4) and (4.6), respectively. The experiment employs a template image of Einstein and a ground truth image of Newton, both rotated in 2D over a uniform polar grid with $L=30$ samples. The reconstructed images within the light-blue rectangle (right panel) show superior performance with the MMSE rotation estimator, with Pearson cross-correlation (PCC) provided for each reconstructed image. The MAP and MMSE reconstructions are nearly identical at high SNR ($\sigma\to 0$), as predicted by Proposition 2.1. At very low SNR ($\sigma\to\infty$), the “Einstein from Noise” effect appears, where the estimator resembles the template image of Einstein rather than the underlying truth of Newton. In intermediate SNR ranges, using the MMSE estimator in the iterative step clearly outperforms the MAP estimator.
@@
Figure 5: Comparison of 3D structure reconstruction in cryo-ET subtomograms averaging using the MMSE and MAP rotation estimators. Iterative structure reconstruction procedures with the MMSE estimator ($\hat{g}_{\text{MMSE}}$) and MAP estimator ($\hat{g}_{\text{MAP}}$) are defined in (4.4) and (4.6), respectively. At high SNR levels, a low-resolution structure emerges due to finite grid sampling of the rotation group $\mathsf{SO}(3)$, effectively acting as a low-pass filter. The 3D reconstruction using the MMSE estimator consistently outperforms the reconstruction using the MAP estimator. For high SNR conditions (i.e., $\sigma\to 0$), both estimators yield similar 3D structures, as expected from Proposition 2.1. The SNR values used for the panels (from right to left) are $10^{-2}$, $2\times 10^{-3}$, $7\times 10^{-4}$, and $2\times 10^{-4}$. The boxes highlighted in purple resemble the true input structure (Ribosome-S80 [59]), while the orange-highlighted boxes are more similar to the initial template (beta-galactosidase [6]), illustrating the “Einstein from Noise” phenomenon. Notably, at very low SNRs, the “Einstein from Noise” effect is evident with the MAP estimator but not with the MMSE estimator. In the “Einstein from Noise” regime, high-frequency details remain unfiltered, revealing finer structural features, in contrast to the suppressed high-frequency information in the accurate 3D reconstruction.

### v2

@@ 4 Bayesian orientation estimation as part of the volume reconstruction problem
Thus far, we have introduced the MLE, MAP and MMSE orientation estimators for estimating the rotation between a single noisy observation $y$ and a reference volume V{V}, as defined in (2.1). An intriguing question that arises is how these estimators can be utilized and influence performance within a 3D volume reconstruction process, which constitutes the main computational challenge in cryo-EM and cryo-ET.
@@
Structure reconstruction typically follows two primary approaches: hard-assignment or soft-assignment methods, which are generally implemented through iterative refinement. In the hard-assignment approach, each observation is assigned a single orientation based on the highest correlation, and the 3D structure is then reconstructed given the rotations. In contrast, the soft-assignment method assigns probabilities across all possible orientations for each observation, enabling the 3D structure to be recovered as a weighted average of the observations, with weights determined by these probabilities. The iterative application of the soft-assignment procedure aligns with the EM algorithm dempster1977maximum ; sigworth2010introduction , which serves as the core computational method in modern cryo-EM scheres2012relion ; punjani2017cryosparc .
@@
In the following, we demonstrate that incorporating the MMSE orientation estimator into the volume reconstruction process in the cryo-ET model (without projections) resembles the expectation-maximization (EM) algorithm, as it accounts for the full distribution of possible outcomes. In contrast, substituting the MLE estimator into the algorithm operates more like a hard-assignment reconstruction method, focusing exclusively on the most likely outcome. This distinction underscores the broader applicability and flexibility of the Bayes estimator in capturing uncertainty and delivering more accurate estimates for structure reconstruction.
@@ 4.1 Connection to the EM algorithm in volume reconstruction without projection
Unlike the previous orientation estimation model (2.1), we consider the following simplified model for volume reconstruction (see also singer2018mathematics ; bandeira2023estimation ; fan2024maximum ). We observe $M$ i.i.d samples taking the form
$y_{i}=g_{i}\circ V+\varepsilon_{i},\quad\quad i=0,\ldots,M-1,$ (4.1)
where $V$ is 3D volume structure of interest, $\{g_{i}\}_{i=0}^{M-1}\in\mathsf{SO}(3)$ are unknown latent variables following i.i.d. uniform distribution, satisfying $g_{i}\circ V(x)\equiv V(g_{i}^{-1}x)$, and $\{\varepsilon_{i}\}_{i=0}^{M-1}$ are i.i.d isotropic Gaussian noise with variance $\sigma^{2}$. For the case where $V$ represents a 2D image, $\{g_{i}\}_{i=0}^{M-1}\in\mathsf{SO}(2)$ corresponds to in-plane rotations, and the model is used for image recovery (see more in Section 4.2ma2019heterogeneous . The goal is to recover $V$ from the observations $\{y_{i}\}_{i=0}^{M-1}$, treating the rotations $\{g_{i}\}_{i=0}^{M-1}$ as latent variables.
@@
The most common method for solving this reconstruction problem is the EM algorithm, which applies soft assignment iteratively, as outlined in Algorithm 2. In each iteration, the algorithm uses the volume estimate from the previous iteration, denoted as $\hat{V}^{(t)}$, to update the volume estimate $\hat{V}^{(t+1)}$, based on the observations $\left\{y_{i}\right\}_{i=0}^{M-1}$. The following proposition illustrates the relationship between the volume structure update rule at iteration $t+1$, and the MMSE orientation estimator introduced in (2.11).
@@
    Compute for every $0\leq i\leq M-1$:
    $p_{i}^{(t)}(g)=\frac{\exp\Big(-\left\|y_{i}-\left(g\circ\hat{V}^{(t)}\right)\right\|^{2}/2\sigma^{2}\Big)}{\int_{g\in\mathsf{SO}(3)}\exp\Big(-\left\|y_{i}-\left(g\circ\hat{V}^{(t)}\right)\right\|^{2}/2\sigma^{2}\Big)\mathrm{d}g},$ (4.2)
@@
    $\hat{V}^{(t+1)}=$ $\underset{V}{\operatorname*{arg\,min}}\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left\|y_{i}-\left(g\circ V\right)\right\|^{2}\mathrm{d}g$
    $=$ $\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)(g^{-1}\circ y_{i})\mathrm{d}g.$ (4.3)
@@
$\hat{V}^{(t+1)}=\frac{1}{M}\sum_{i=0}^{M-1}\hat{\mathfrak{g}}_{\mathsf{MMSE},i,t}\circ y_{i},$ (4.4)
@@
where $\hat{\mathfrak{g}}_{\mathsf{MMSE},i,t}$ denotes the MMSE rotation operator defined as the posterier mean of $g^{-1}$ with respect to the uniform prior, given observation $y_{i}$ and underlying structure $\hat{V}^{(t)}$, using the same posterior weight as in (2.11):
$\hat{\mathfrak{g}}_{\mathsf{MMSE},i,t}:=$ $\mathbb{E}[g^{-1}|y_{i},\hat{V}^{(t)}]$
$\mathrel{\text{``=''}}$ $\int_{g\in\mathsf{SO}(3)}\ p_{i}^{(t)}(g)g^{-1}\mathrm{d}g,$ (4.5)
where the quoted equality emphasizes that the expectation (or integral) is interpreted as acting on the operator, in contrast to the previous $3\times 3$ posterior matrix mean in (2.11).
@@
The proof of this proposition is presented in Appendix F. In words, the proposition shows that the update rule for the volume structure estimation at iteration $t+1$, given the volume $\hat{V}^{(t)}$, is equivalent to aligning each observation $\left\{y_{i}\right\}_{i=0}^{M-1}$ using the associated MMSE operator $\hat{\mathfrak{g}}_{\text{MMSE},i,t}$, computed based on the reference volume $\hat{V}^{(t)}$, and then averaging the aligned observations (4.4). Thus, the MMSE rotation operator is a key ingredient in the EM algorithm for volume reconstruction.
Remark 4.2.
The MMSE rotation “operator” introduced above is closely related to the MMSE rotation “estimator”, namely the orthogonal Procrustes projection of the matrix-based posterior mean $\hat{g}_{\mathsf{relax}}$ discussed in (2.12). In the matrix formulation, the expectation produces a $3\times 3$ posterior matrix mean. In contrast, in the operator formulation, the expectation acts directly on the data $y_{i}$, effectively performing a posterior-weighted average over $g^{-1}$. While the two formulations differ in representation, the underlying principle is the same: both correspond to a posterior-weighted aggregation of rotations according to the posterior distribution.
Moreover, if the volume is expressed in a suitable basis, such as the spherical harmonics basis, the operator has a corresponding matrix representation (e.g., a block-diagonal matrix of rotation matrices). In this case, the MMSE operator can also be written as a posterior mean of such block-diagonal rotation matrix.
Despite the subtle differences arising from rotation estimation versus structure reconstruction, the previously introduced MMSE rotation estimator (2.11) can still serve as a practical tool for assessing the performance of soft assignments. Concretely, we take the $3\times 3$ posterior rotation matrix mean as in (2.11),
$\hat{g}_{\mathsf{relax},i,t}:=\int_{g\in\mathsf{SO}(3)}\ p_{i}^{(t)}(g)g\mathrm{d}g,$
perform an Orthogonal Procrustes projection to enforce a valid rotation, and then take the inverse to replace $\hat{\mathfrak{g}}_{\mathsf{MMSE},i,t}$. In this way, the MMSE rotation estimator provides a concrete approximation to the soft assignment in structure reconstruction. In numerical experiments, it is compared with the hard assignment, demonstrating superior performance relative to the maximum-likelihood estimator, as discussed below.
The MLE estimator as part of volume reconstruction.
@@
In contrast to the soft assignment procedure, if we replace the MMSE operator with the corresponding MLE operator, we obtain the structure reconstruction algorithm by applying hard assignment iteratively. To be more specific, the MLE operator applied to the observations $\left\{y_{i}\right\}_{i=0}^{M-1}$ can be viewed as a hard assignment among all possible rotations. In practice, this procedure involves making a hard decision where a single rotation is selected from the rotation grid according to the closest alignment. In the $(t+1)$-th iteration, similarly to (4.4), the hard-assignment process can be expressed as follows:
$\hat{V}^{(t+1)}=\frac{1}{M}\sum_{i=0}^{M-1}\hat{\mathfrak{g}}_{\mathsf{MLE},i,t}\circ y_{i},$ (4.6)
where $\hat{\mathfrak{g}}_{\mathsf{MLE},i,t}$ is defined by
$\hat{\mathfrak{g}}_{\mathsf{MLE},i,t}=$ $\operatorname*{arg\,max}_{g^{-1}\in\mathsf{SO}(3)}{\mathbb{P}}(g|y_{i},\hat{V}^{(t)})$ (4.7)
$=$ $\operatorname*{arg\,max}_{g^{-1}\in\mathsf{SO}(3)}p_{i}^{(t)}(g)=\operatorname*{arg\,min}_{g^{-1}\in\mathsf{SO}(3)}\|y_{i}-(g\circ\hat{V}^{(t)})\|^{2}.$ (4.8)
In other words, $\hat{\mathfrak{g}}_{\mathsf{MLE},i,t}=(\hat{g}_{\mathsf{MLE},i,t})^{-1}$ where
$\hat{g}_{\mathsf{MLE},i,t}:=\operatorname*{arg\,min}_{g\in\mathsf{SO}(3)}\|y_{i}-(g\circ\hat{V}^{(t)})\|^{2}.$
Hence, the hard assignment can be viewed as using the exact inverse of the MLE estimator introduced earlier.
@@ 4.2 Empirical results for volume reconstruction and the “Einstein from Noise” phenomenon
We demonstrate empirically volume reconstruction by applying the MLE estimator $\hat{g}_{\mathsf{MLE}}$ and the MMSE estimator $\hat{g}_{\mathsf{MMSE}}$ as part of the reconstruction problem, as specified in Section 4.1.
@@
We demonstrate the reconstruction processes, which integrates MLE and MMSE rotation estimators as intermediate steps, using the two examples: 2D image recovery (Figure 4) and 3D volume reconstruction without projection (Figure 5). The iterative reconstruction process, as outlined in Algorithm 2, was performed until convergence (i.e., when the relative difference between consecutive iterations fell below a predefined threshold of $10^{-3}$ or until reaching a maximum of 100 iterations.
@@
Figures 4 and 5 were generated using slightly different methods. For the 2D experiment presented in Figure 4, we used polar coordinates, while in the 3D experiment, shown in Figure 5, we used a standard Cartesian basis. The primary difference lies in the interpolation required for producing Figure 5, which utilizes cubic interpolation for each observation based on the estimated rotation. This introduces certain “quantization” errors. Additionally, as detailed in Section 2, the 3D reconstruction process, presented in Figure 5, requires a “rounding” step which amounts to solving the Orthogonal Procrustes procedure. In Figure 4, the true and template structures were generated in a polar representation with $d=300$ radial points and $L=30$ polar angle points. The reconstruction process was performed with $M=5\times 10^{4}$ observations. The additive noise was added in the polar representation. In Figure 5, $M=3000$ observations were used, with the rotation group $\mathsf{SO}(3)$ grid size of $L=300$, and a volume size of $32\times 32\times 32$.
@@
Figure 4: Comparison of 2D image recovery using the MMSE and the MLE rotation estimators. Iterative image recovery procedures with the MMSE estimator ($\hat{g}_{\text{MMSE}}$) and MLE estimator ($\hat{g}_{\text{MLE}}$) are defined in (4.4) and (4.6), respectively. The experiment employs a template image of Einstein and a ground truth image of Newton, both rotated in 2D over a uniform polar grid with $L=30$ samples. The reconstructed images within the dark-orange rectangle (right panel) show superior performance with the MMSE rotation estimator, with Pearson cross-correlation (PCC) provided for each reconstructed image. The MLE and MMSE reconstructions are nearly identical at high SNR ($\sigma\to 0$), as predicted by Proposition 2.2. The SNR values used for the panels (from right to left) are $10^{-2}$, $4\times 10^{-3}$, $2\times 10^{-3}$, $7\times 10^{-4}$, and $2\times 10^{-4}$. At very low SNR ($\sigma\to\infty$), the “Einstein from Noise” effect appears, where the estimator resembles the template image of Einstein rather than the underlying truth of Newton. In intermediate SNR ranges, using the MMSE estimator in the iterative step clearly outperforms the MLE estimator.
@@
Figure 5: Comparison of 3D structure reconstruction in cryo-ET subtomograms averaging using the MMSE and MLE rotation estimators. Iterative structure reconstruction procedures with the MMSE estimator ($\hat{g}_{\text{MMSE}}$) and MLE estimator ($\hat{g}_{\text{MLE}}$) are defined in (4.4) and (4.6), respectively. At high SNR levels, a low-resolution structure emerges due to finite grid sampling of the rotation group $\mathsf{SO}(3)$, effectively acting as a low-pass filter. The 3D reconstruction using the MMSE estimator consistently outperforms the reconstruction using the MLE estimator. For high SNR conditions (i.e., $\sigma\to 0$), both estimators yield similar 3D structures, as expected from Proposition 2.2. The SNR values used for the panels (from right to left) are $10^{-2}$, $2\times 10^{-3}$, $7\times 10^{-4}$, and $2\times 10^{-4}$, with a volume size of $32\times 32\times 32$. The boxes highlighted in purple resemble the true input structure (Ribosome-S80 wong2014cryo ), while the orange-highlighted boxes are more similar to the initial template (beta-galactosidase bartesaghi2014structure ), illustrating the “Einstein from Noise” phenomenon. Notably, at very low SNRs, the “Einstein from Noise” effect is evident with the MLE estimator but not with the MMSE estimator.